# Modular Spiking Neural Networks

### v1

@@ 1 Introduction
Current artificial intelligence systems inspired by the brain have shown remarkable capabilities in most domains of human interest. However, the predominant flavor of neural networks is significantly inefficient, both in terms of energy and information when compared to natural neural networks (also known as brains) (Stiefel and Coggan, 2023; Lake et al., 2017). Several alternative artificial neural models have been proposed, with spiking neural networks (SNNs) being the most popular among them due to the possibility of efficient hardware implementations (Stanojevic et al., 2024) and the collective hunch that effective plasticity mechanisms may open the door to efficient learning as in natural neural networks (Lake et al., 2017).
@@ 3.2 Performance
Our testbed consists of three different network architectures, 25 network instantiations per architecture and 5 runs per instantiation for a span of 10 seconds. Simulations are given a timeout of 60 seconds, after which, simulation time is estimated based on the number of iterations completed. Simulations were conducted assuming different interaction times: $1ms$, $2ms$, $5ms$, $10ms$ and $50ms$. An interaction step indicates how much time each model computes before a new input signal is presented to the network and a new output signal is read from the network. Write-in and read-out operations are expensive in CPU’s and even more so in GPU’s. A good balance of write-in and read-out is necessary to get the most out of a SNN. Note that use of the concepts of "interaction times" and "msms" here may be an overstretch, and they are just a bioinspired abstraction for how much computation do we perform between I/O operations.
@@
Brian2 also offers GPU implementations (Alevi et al., 2022; Stimberg et al., 2020), although only the CUDA implementation is actively maintained. Unlike the other simulations, we ran Brian2 CUDA in a Docker container due to incompatibilities with our current setup. Moreover, our benchmark network architectures involve a substantial number of spike propagation operations which are known to significantly slow down the CUDA backend (Alevi et al., 2022); we tested several configurations using one of our network architectures and ran all simulations with the most performant setup we found.
Runtimes and compilation times are shown in figure 4. As expected, the performance of all the simulations increases as the interaction time increases: stopping a simulation to do an interactive step is a costly operation, even when no cross-communication between the CPU and the GPU is necessary. Spark significantly outperforms Brian2 implementations, regardless of the backend. Our framework is up to $\sim 5$ to $\sim 200$ times faster, depending on the interaction steps, than running a single execution of Brian2 C++ on our testbed. Major speedups, of up to 3 orders of magnitude, are observed when compared to the rest of the interaction mock-up simulations. Brian2 CUDA was slower than all CPU-based implementations in our testbed. According to Alevi et al. (2022), our network size ($\sim 10^{4}$ units per pool) may be too small for Brian2CUDA to take advantage and, as mentioned above, spike propagation operations are quite abundant in our simulations. This operation severely impacts CUDA performance and can even make it slower than the C++ backend. Note that Brian2 offers several advanced features for direct manipulation of the compiled code which, in theory, may be used to increase the performance for interactive simulations. However, such modifications may require a deep understanding of Brian2’s implementation, making them unavailable to most SNN practitioners.

### v2

@@ 1 Introduction
Current artificial intelligence systems inspired by the brain have shown remarkable capabilities in most domains of human interest. However, the predominant flavor of neural networks is significantly inefficient, both in terms of energy and information when compared to natural neural networks (also known as brains) (Stiefel and Coggan, 2023; Lake et al., 2017). Several alternative artificial neural models have been proposed, with spiking neural networks (SNNs) being the most popular among them due to the possibility of efficient hardware implementations (Stanojevic et al., 2024; Roy et al., 2025) and the collective hunch that effective plasticity mechanisms may open the door to efficient learning as in natural neural networks (Lake et al., 2017).
@@ 3.2 Performance
Our testbed consists of three different network architectures, 25 network instantiations per architecture and 5 runs per instantiation for a span of 10 seconds. Simulations are given a timeout of 60 seconds, after which, simulation time is estimated based on the number of iterations completed. Simulations were conducted assuming different interaction times: $1ms$, $2ms$, $5ms$, $10ms$, $50ms$ and $100ms$. An interaction step indicates how much time each model computes before a new input signal is presented to the network and a new output signal is read from the network. Write-in and read-out operations are expensive in CPU’s and even more so in GPU’s. A good balance of write-in and read-out is necessary to get the most out of a SNN. Note that use of the concepts of "interaction times" and "msms" here may be an overstretch, and they are just a bioinspired abstraction for how much computation do we packed between I/O operations.
@@
Brian2 also offers GPU implementations (Alevi et al., 2022; Stimberg et al., 2020), although only the CUDA implementation is actively maintained. Moreover, our benchmark network architectures involve a substantial number of spike propagation operations which are known to significantly slow down the CUDA backend (Alevi et al., 2022); we tested several backend configurations using one of our network architectures and ran all simulations with the most performant setup.
Runtimes and compilation times are shown in figure 4. As expected, the performance of all the simulations increases as the interaction time increases: stopping a simulation to do an interactive step is a costly operation, even when no cross-communication between the CPU and the GPU is necessary. Spark significantly outperforms Brian2 implementations, regardless of the backend. Our framework is up to $\sim 5$ to $\sim 350$ times faster, depending on the interaction steps, than running a single execution of Brian2 C++ on our testbed. Major speedups, of up to 3 orders of magnitude, are observed when compared to the rest of the interaction mock-up simulations. Brian2 CUDA was slower than all CPU-based implementations in our testbed. According to Alevi et al. (2022), our network size ($\sim 10^{4}$ units per pool) may be too small for Brian2CUDA to take advantage and, as mentioned above, spike propagation operations are quite abundant in our simulations. This operation severely impacts the CUDA backend performance and can even make it slower than the C++ backend. Note that Brian2 offers several advanced features for direct manipulation of the compiled code which, in theory, may be used to increase the performance for interactive simulations. However, such modifications may require a deep understanding of Brian2’s implementation, making them unavailable to most SNN practitioners.